OpenStudy (anonymous):

Use the distributive property to simplify the expression. -7(8x-3)+1 A. x+10 B.-15x+4 C.-9x-4 D. -55x+21

OpenStudy (anonymous):

I got D.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Something is wrong in the way it is written. :) 7*8 = 56

OpenStudy (anonymous):

None are correct. That simplifies to -56x+22

OpenStudy (anonymous):

no it doesn't @Vandreigan it has to be one of those

OpenStudy (anonymous):

7*8 = 56. 3*7 = 21. 21+1 = 22.

OpenStudy (anonymous):

You must have written the expression wrong. It might be: -7(8x-3) + x I meant.

OpenStudy (anonymous):

yes I'm sorry everyone the expression is -7(8x-3)+x and i got D,is that correct?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

thanks:)